• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 22:32:07 on October 7,2009

    Entry number 293733

    keyword=Compton position scan

    Laser on retes were low after beam returned (500Hz/uA), so the Compton orbit lock position has been changed to IPM1P02B.YPOS=-0.95 (from -0.8) to give laser on rates of 790Hz/uA.

    FIGURE 1